# Global Prediction of COVID-19 Variant Emergence Using Dynamics-Informed Graph Neural Networks

### v2

@@ V-C Implementation
The main pipeline depicted in Figure 5, was implemented in Python. The dynamics-based model and dynamics-based delay prediction models were implemented using MATLAB, following a similar overall pipeline. The implementation was carried out on a machine equipped with a 32-core Intel(R) Xeon(R) Gold 5218 CPU running at 2.3 GHz and 64 GB of RAM. Focusing on the example of variant 23F.Omicron at a single timestep, which contains the majority of the data, and the heaviest model (T-GCN), the average runtimes for different components were approximately 32s for data pre-processing, 20s for training, and 0.01 seconds for inference. The entire retrospective pipeline takes approximately 2 hours to complete. Notably, the most time-consuming aspect of the retrospective pipeline is the data pre-processing step, which could potentially be optimized further through memoization.
@@ VI Conclusion
We addressed the challenges of predicting variant emergence and delay across countries. The derivation of variant dynamics provided a theoretical foundation, which was used to engineer relevant features as well as a novel dynamics-based model. We demonstrated that our dynamics-informed GNN approach outperformed all other methods. Through comprehensive experiments and analysis, we demonstrated the effectiveness of our design choices, providing valuable tools for understanding and predicting the intricate relationships and connectivity patterns between nations. Furthermore, as we are the first to address the proposed problem, we provided a comprehensive benchmark and make our full pipeline available in an effort to facilitate research in the field.

### v3

@@ V-C Implementation
The pipeline depicted in Figure 5, was implemented in Python. The implementation was carried out on a machine equipped with a 32-core Intel(R) Xeon(R) Gold 5218 CPU running at 2.3 GHz and 64 GB of RAM. Focusing on the example of variant 23F.Omicron at a single timestep, which contains the majority of the data, and the heaviest model (FA-GCN), the average runtimes for different components were approximately 32 seconds for data pre-processing, 20 seconds for training, and 0.01 seconds for inference. The entire retrospective pipeline takes approximately 2 hours to complete. Notably, the most time-consuming aspect of the pipeline is the data pre-processing step, which could potentially be optimized further through memorization.
V-D Limitations
One limitation of our approach in addressing the problem of Delay Prediction lies in handling situations where a variant might not even become dominant in a given country. To tackle this issue, we define the following problem:
Problem 2 (Dominance Prediction).
Given the prevalence (proportion) of a variant in regions $A_{1},A_{2},\dots$, predict if the variant will ever reach a proportion $\theta$ in region $B$.
In future work, we plan to address this issue by adopting a two-step approach. First, we solve Dominance Prediction, which aims to predict whether a variant will become dominant in a given region. If the variant is predicted to become dominant, we can then proceed to solve Delay Prediction.
@@ VI Conclusion
We addressed the challenges of predicting variant delay across countries. The derivation of variant dynamics provided a theoretical foundation, which was used to engineer relevant features as well as a novel baseline dynamics-based model. We demonstrated that our dynamics feature-augmented GNN approach outperformed all other methods. Through comprehensive experiments and analysis, we demonstrated the effectiveness of our design choices, providing valuable tools for understanding and predicting the intricate relationships and connectivity patterns between nations. Furthermore, as we are the first to address the proposed problem, we provided a comprehensive benchmark and made our full pipeline available in an effort to facilitate research in the field.
Acknowledgements
This work was supported by the Centers for Disease Control and Prevention and the National Science Foundation under the awards no. 2135784, 2223933, and 2333494.